Suppose, in a two sector model comprising consumers and firms, that the individuals receive the following payments for their provision of goods and services: wages K60000, interest K30000, rent K10000 and profits k8000. Consumption spending is K101000, planned investment expenditure is K20000 and actual investment expenditure is K7000.
a) What is the equilibrium output? b) What is the level of saving in this economy; clearly show how it’s calculated. c) What is the relationship of saving and investment? d) What is a leakage from the circular flow of income? e) Why is investment spending viewed as an injection into the circular flow of income? 10 marks
1
Expert's answer
2018-03-05T08:35:07-0500
A) GDP = W + R + P + R = K60000 + K30000 + K10000 +K8000 = K108000
B) Level of saving in this economy = (K108000 – K101000) = K7000
C) Saving and investment are equal only in equilibrium; they are unequal under
conditions of disequilibrium.
D) In a closed circular income stream, money flows continuously from firms to
households. Households spend all of their money on goods and businesses spend
all of their money on labor and expansion. The basic model of the circular flow of
income ignores common consumer actions that take money out of the circular of
income, or leakage. For instance, most households save money and some
companies retain earnings.
E) The circular flow will adjust following new injections into it or new withdrawals
from it. An injection of new spending will increase the flow. A net injection relates
to the overall effect of injections in relation to withdrawals following a change in
an economic variable.